AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 438     Entry time: Thu Aug 31 08:36:34 2023
Author: CB 
Subject: NEG activation  

Found CARME at 50C, a bit cooler than expected, probably due to invasive tent changes yesterday to mount missing wires. Increased heaters temperature to 300C

 

9:35 Started D-2000 #1,3 (behind the DSSD). T=400, tramp 30, thold 90

11:20 Set D-2000 1,3 to 200C. Turned on D-2000 2,4 as above 

13:15 Set D-2000 2,4 to 200C. Turned on UHV1400 set 1 at 250W, tramp 30, thold infinity (will ramp down manually as required)

13:20 No appreciable effects on DSSD temperature. Turned on set 2 as well, which should be mounted close nearby.

13:45 Both sets reached 250W. No effects on detector temperature. Increased to 300W. Vacuum gets slightly worse, but recovers quickly. Not worth further increasing power. Leaving for 90 minutes.

15:20 Strain relief TC at 75C quite hotter than detector TC at 35C. Might be normal. Vacuum division IE514 must have been degassed remotely? Decreased set 1,2 to 75W. Turned on sets 3,4 to 250W as above. 

15:45 Sets 3, 4 now at 250W. Detector TC heating quickly at 13C/h but nowhere near dangerous temperatures yet. Decreased sets 1,2 to 50W. Leaving for 90 minutes.

17:15 Sets 3, 4 to 50W. Slightly worrying temperature difference between detectors and strain relief around 65C. Different materials and different positions could explain this.

17:40 Oddly, detector temperature now going up sharply (temperature diff. down to 40C) since the uhv1400 power was decreased. Cannot think of s logical reason,  it may just be a coincidence and heat transfer is catching up.

 

Leaving overnight. Will turn on the last two groups of UHV1400 tomorrow morning.

ELOG V3.1.4-unknown